I don't know how to solve this equation.
Use factoring to solve each equation for real values of the variable.
Please help me solve this problem: 6p + p^1/2 - 1 = 0

Answer by checkley77(12844)

6p+p^1/2-1=0
(3SQRTP-1)(2SQRTP+1)=0
3SQRTP-1=0
3SQRTP=1 [SQUARE BOTH SIDES.]
9P=1
P=1/9 ANSWER.
2SQRTP+1=0
2SQRTP=-1 [SQUARE BOTH SIDES]
4P=1
P=1/4 ANSWER.